# Advancing Averaged Primer Vector Theory with Bang-Bang Control and Eclipsing

### v1

@@ Abstract
Low-thrust, many-revolution spacecraft trajectories are increasingly required for mission design due to the efficiency and reliability of electric propulsion technology. Primer vector theory using averaged dynamics is well suited for such applications, but is difficult to implement in a way that maintains both optimality and computational efficiency. An improved model is presented that combines advances from several past works into a general and practical formulation for minimum-fuel, perturbed Keplerian dynamics. The model maintains computational efficiency of dynamics averaging with optimal handling of the eclipsing constraint and bang-bang control through the use of the Leibniz integral rule for multi-arc averaging. A subtle, but important singularity arising from the averaged eclipsing constraint is identified and fixed. A maximum number of six switching function roots per revolution is established within the averaged dynamics. This new theoretical insight provides a practical upper-bound on the number of thrusting arcs required for any low-thrust optimization problem. Variational equations are provided for fast and accurate calculation of the state transition matrix for use in targeting and optimization. The dynamics include generic two-body perturbations and an expanded state to allow for sensitivity calculations with respect to launch date and flight time. A 48-revolution GTO to GEO transfer is used to directly compare optimal averaged and unaveraged trajectories. The capabilities of averaged dynamics are then demonstrated with an optimal 486-revolution GTO to GEO minimum fuel transfer.
[…]
@@ 1 Introduction
The increasing use of low-thrust, high-efficiency propulsion systems is making long flight-time, many-revolution trajectories (on the order of hundreds of revolutions) more relevant.33footnotemark: 300footnotetext: 33footnotemark: 3A preliminary version of this paper was presented at the 2024 AIAA SciTech Forum [1]. While it can often be sufficient to use a simple control law for these many-revolution trajectories [2, 3, 4], optimization can provide significant savings in fuel and/or time of flight, and potentially enable new missions or extend the capability of existing missions. In the mission design process, the need for repeated trajectory optimization (e.g. trade studies) requires optimization methods that are efficient and robust.
@@
The indirect approach using primer vector theory [5] is a popular choice for handling low-thrust, many revolution spacecraft trajectories [6]. This method transforms the optimization problem into a nonlinear two-point boundary value problem by introducing Lagrange multipliers (or co-states) to solve for the optimal control law throughout the trajectory. The problem is reduced to finding the initial values of the co-states. While the indirect method benefits from low dimensionality, the problems are notoriously hard due to high numerical sensitivity and complexity associated to incorporating constraints. In 1965 and 1967, Edelbaum used optimal control to analytically solve for low-thrust transfers between nearby elliptical orbits [7, 8]. His approach was extended to the general case of transfers between arbitrary elliptical orbits by Marec and Vinh [9]. Kechichian improved on these methods by using equinoctial elements to reduce the number of singularities [10]. More recent works have successfully applied the indirect method to many revolution trajectories using a variety of approaches, such as optimal handling of bang-bang control with switching detection [11, 12], a smooth approximation of bang-bang control [13, 14], and interior point constraints [11, 15].
@@
In 1973, Edelbaum et al. introduced the use of dynamics averaging in optimal control for minimum time problems [16], which increased optimization speed and robustness. This averaging method was then used in the SESPOT software developed by him, Sackett, and Malchow [17]. Marec and Vinh extended optimal control with dynamics averaging to the minimum fuel problem [18]. Kluever and Conway [19] build on SESPOT with a model that calculates the control profile by discretizing the co-state time histories and solving the resulting NLP, although the model does not include variations in right ascension of the ascending node or argument of periapse. Geffroy and Epenoy [20] incorporated constraints on thrust direction and placed optimal control with dynamics averaging on more sound mathematical footing using Chaplais’ theory of averaging [21]. While the effect of Earth’s shadow had been included in previous averaging models, it was incorporated in a technically optimal manner for the first time by Mazzini [22]. Tarzi [23] incorporated solar radiation pressure into the model and introduced a multi-arc approach to averaging for bang-bang control, but did not account for the Earth’s shadow in an optimal way. D’argent [24, 25] introduces a model that incorporates second-order averaging and explores the transformation between averaged and unaveraged solutions, but does not include eclipsing. Wu et al. [26] develop a method to analytically calculate the roots of the switching function in the averaging context and use a neural network for initial guess generation, but does not include eclipsing in an optimal manner or in examples. Many of these more recent averaging methods are compared by Petropolous [27]. The objective of this current work is to combine and extend the advances for the averaged dynamics case into a general and practical formulation for minimum fuel problems. The new model includes optimal handling of the eclipsing constraint, multi-arc averaging with novel application of the Leibniz integral rule, inclusion of generic dynamics perturbations, and variational equations for averaged dynamics for the first time.
@@
Dynamics averaging removes the fast periodic changes in the states and co-states, leaving only the slowly changing secular motion. The resulting averaged trajectory will approximate the unaveraged trajectory to within , for integration times of up to [28], where is typically a small parameter defined in terms of perturbation amplitude (control or dynamics). A numerical integrator can then take much larger time steps, which both reduces the numerical sensitivity to the initial co-states and reduces computation time to solve the resulting system of ordinary differential equations.
@@
In practice, dynamics averaging is done by taking the Hamiltonian (with the embedded primer vector feedback law) and integrating over one period in time to find the averaged Hamiltonian. Derivatives needed for equations of motion (and the variational equations) are then calculated using the averaged Hamiltonian. Following the methods of Kechichian [29] and Edelbaum et al. [16], the averaging integrals are approximated using a weighted sum with Gaussian Quadrature because they rarely can be calculated analytically. For the minimum fuel problem with bang-bang control, a significant number of quadrature points are required to accurately capture the thrusting and coasting proportions of the averaging period. The multi-arc averaging approach introduced by Tarzi [23] for bang-bang control separates the integral over the Hamiltonian into different thrusting or ballistic “arcs”. The multi-arc approach reduces the number of quadrature points required to achieve a specific level of accuracy at the cost of root finding in each dynamics function call.
@@
One of the most important constraints for low-thrust trajectory design is the effect of the Earth’s shadow, which typically requires a spacecraft with an electric propulsion engine to stop thrusting. Incorporating this constraint in a way that maintains optimality is not simple in general, and especially difficult when using dynamics averaging. In a proper optimal control derivation, the eclipsing constraint is directly incorporated into the dynamics so that the control law is optimal with respect to the constraint [30, 31]. If, instead, the constraint is simply enforced in flight, using a thrust profile that did not consider eclipsing, then the control law is technically no longer optimal. Mazzini describes the importance of the dynamics-based control constraint with multiple examples, showing how it can provide significant fuel savings depending on the Sun position at launch [22]. When the constraint is properly incorporated, the co-states exhibit a discrete “jump” at eclipse entrance and exit. In averaged dynamics, these discrete jumps in a co-state variable are not easily incorporated. Mazzini [22] develops a method to directly calculate the discrete jumps by integrating approximate co-state dynamics through a transitionary layer for eclipsing. The averaged co-state jumps are then calculated by applying the averaging integral to co-state equations of motion that include a Dirac delta function. The resulting form of the averaged co-state jumps is sufficient for propagation of an averaged trajectory, but does not address calculation of the variational equations. The model introduced in this current work incorporates the co-state jumps into the averaged dynamics using the Leibniz integral rule applied to multi-arc averaging. This new approach uses no approximations and enables straightforward calculation of the variational equations.
@@
In the current work, the model includes the multi-arc averaging approach introduced by Tarzi [23] along with a novel application of the Leibniz integral rule to automatically incorporate the averaged co-state jumps. The new approach is simple, efficient, optimal with respect to the eclipsing constraint, and allows for straight-forward calculation of the variational equations. The averaged co-state jumps, however, approach a singularity as eclipsing arcs approach zero length. This singularity, also identified in Mazzini’s model, can significantly slow down or prevent propagation with a variable step integrator and result in accumulation of error. The new averaging model removes this singularity through a simple redefinition of the eclipsing constraint.
@@ 2 Optimal Control
Similar to prior works on this many-revolution low-thrust problem [32, 14, 33], modified equinoctial orbital elements are used to reduce the number of singularities compared to Keplerian orbital elements [34]. Current time, , and overall time of flight, , are included in the state variables:
@@
In order to incorporate the effects of the Earth’s shadow directly into the dynamics, the conventional geometric “eclipsing function” is used [30]. The eclipsing function is a smooth function of the geocentric satellite position, the geocentric Sun position, the radius of the Sun, the radius of the Earth, and indirectly the state variable time. A conical shadow model is used, where the engine fully shuts off in umbra, penumbra, and antumbra (see Fig. 1)[35]. The eclipsing function is written as
@@
The spacecraft thrust is defined in the radial-transverse-normal (RTN) reference frame. The controllable variables are the thrust direction, , and a thrust modulating parameter, , with maximum and minimum thrusts given as and , respectively. The function is incorporated directly into the dynamics. The equations of motion are defined:
@@
where is an arbitrary perturbation, is the gravitation acceleration at Earth’s surface, and is the specific impulse. The matrices A, B contain the state dynamics equations for modified equinoctial orbital elements written in the RTN frame. They are derived by Kechichian [29] and shown in the appendix. The arbitrary perturbation, , and the thrust unit vector, , are defined in the RTN frame, and must be a smooth function of state variables. This perturbation must be small enough to not significantly change the mean orbital elements over one period for the averaging assumptions to hold. Examples include oblateness terms for the primary or simple solar radiation or drag models.
@@
The minimum fuel cost function [36] used in this model is defined as:
@@
where is the co-state vector for the 6 orbital elements and two time related variables, and is the co-state for the mass. The co-state dynamics are then calculated as:
@@
To derive the optimal control law, the Hamiltonian is minimized with respect to the controls, thrust direction and thrust magnitude [37]. For the direction, the thrust vector is anti-aligned with B [5]:
where is the so-called switching function:
Minimizing the Hamiltonian, Eq. 13, with respect to leads to so-called “bang-bang” control:
@@
This optimal is then substituted back into the Hamiltonian (Eq. 13). The derived optimal control law is now optimal with respect to the eclipsing constraint.
As a result of incorporating the constraint directly into the dynamics in Eq. 8, the co-states naturally exhibit a jump at shadow entrance and exit due to Weierstrass-Erdmann corner conditions [11, 38]. These co-state jumps can be difficult to incorporate into averaged dynamics [22]. This current model smoothly incorporates the jumps using a novel application of the Leibniz integral rule with multi-arc averaging. Next, the averaged dynamics is introduced.
@@ 3.1 Basic method
The averaging approach used in this work follows that of Edelbaum et al. [16] and Kechichian [29]. The process starts with the unaveraged Hamiltonian including the embedded optimal control. Next, the average is performed over one period in time, and this averaged Hamiltonian is used to calculate the equations of motion for the averaged states and co-states. It should be noted that the optimal control does not need to be re-derived after averaging the Hamiltonian, because there are no direct constraints on the control [20]. The averaging integral is converted from time to the fast moving true longitude, L. All other variables are held constant within the averaging integral. This process removes the fast periodic changes in the state and co-state, leaving only the secular motion. Gaussian quadrature [39] is used to approximate the integral and all other equations derived from it. The time averaged Hamiltonian, , over one period is defined here and rearranged slightly:
where a new term, , is defined for convenience and is only a function of the state variables, not the co-states. The state and co-state dynamics are calculated using the chain rule:
@@
where the weights, , and node points, , are calculated from Legendre polynomials [39]. The dynamics equations are then calculated using quadrature:
The number of quadrature points required, n, to make the approximation sufficiently accurate depends on the complexity of the unaveraged dynamics within the given period of averaging. If there are sections of the given period with and without thrusting, then n must be large enough to accurately capture the proportion and relative locations of the two. This common [20, 22], but brute force approach to averaging can sometimes require such large n that the computational savings of averaging are significantly reduced. Alternatively, a multi-arc approach improves the accuracy of approximation without requiring excessively large n.

### v3

@@ Abstract
Primer vector theory using averaged dynamics is well suited for optimizing low-thrust, many-revolution spacecraft trajectories, but is difficult to implement in a way that maintains both optimality and computational efficiency. An improved model is presented that combines advances from several past works into a general and practical formulation for minimum-fuel, perturbed Keplerian dynamics. The model maintains computational efficiency of dynamics averaging with optimal handling of the eclipsing constraint and bang-bang control through the use of the Leibniz integral rule for multi-arc averaging. A subtle, but important singularity arising from the averaged eclipsing constraint is identified and fixed. A maximum number of six switching function roots per revolution is established within the averaged dynamics. This new theoretical insight provides a practical upper-bound on the number of thrusting arcs required for any low-thrust optimization problem. Variational equations are provided for fast and accurate calculation of the state transition matrix for use in targeting and optimization. The dynamics include generic two-body perturbations and an expanded state to allow for sensitivity calculations with respect to launch date and flight time. The new model is illustrated on a GTO to GEO transfer, including up to 486 revolutions.
@@ 1 Introduction
The increasing use of low-thrust, high-efficiency propulsion systems is making long flight-time, many-revolution trajectories (on the order of hundreds of revolutions) more relevant.33footnotemark: 3 22footnotetext: 33footnotemark: 3A preliminary version of this paper was presented at the Jan 2024 AIAA SciTech Forum in Orlando, FL (paper number AIAA 2024-1279 [lifset]) While it can often be sufficient to use a simple control law for these many-revolution trajectories [petropoulos2004low, PPE, wu2021rapid], optimization can provide significant savings in fuel and/or time of flight, and potentially enable new missions or extend the capability of existing missions. In the mission design process, the need for repeated trajectory optimization (e.g. trade studies) requires optimization methods that are efficient and robust.
@@
The indirect approach using primer vector theory [lawden_1963] is a popular choice for handling low-thrust, many revolution spacecraft trajectories [surveypaper]. This method transforms the optimization problem into a nonlinear two-point boundary value problem by introducing Lagrange multipliers (or costates) to solve for the optimal control law throughout the trajectory. The problem is reduced to finding the initial values of the costates. While the indirect method benefits from low dimensionality, the problems are notoriously hard due to high numerical sensitivity and complexity associated to incorporating constraints. In 1965 and 1967, Edelbaum used optimal control to analytically solve for low-thrust transfers between nearby elliptical orbits [edelbaum1965, edelbaum1967general]. His approach was extended to the general case of transfers between arbitrary elliptical orbits by Marec and Vinh [marec1977optimal]. Kechichian improved on these methods by using equinoctial elements to reduce the number of singularities [kechichian1997trajectory]. More recent works have successfully applied the indirect method to many revolution trajectories using a variety of approaches, such as optimal handling of bang-bang control with switching detection [wangtopputo, russell2007primer], a smooth approximation of bang-bang control [TAHERI2020151, taheri], and interior point constraints [wangtopputo, PONTANI2022549, cerf2018].
@@
In 1973, Edelbaum et al. introduced the use of dynamics averaging in optimal control for minimum time problems [edelbaum], which increased optimization speed and robustness. This averaging method was then used in the SEPSPOT software developed by him, Sackett, and Malchow [sackett1975solar]. Marec and Vinh extended optimal control with dynamics averaging to the minimum fuel problem [marec]. Kluever [kluever] builds on SESPOT with a model that calculates the control profile by discretizing the costate time histories and solving the resulting nonlinear problem, although the model does not include variations in right ascension of the ascending node or argument of periapse. Geffroy and Epenoy [geffroy] incorporated constraints on thrust direction and placed optimal control with dynamics averaging on more sound mathematical footing using Chaplais’ theory of averaging [chaplais]. While the effect of Earth’s shadow had been included in previous averaging models, it was incorporated in a technically optimal manner for the first time by Dargent [dargent] using a smoothed model for the constraint (Dargent also investigates the transformation between averaged and unaveraged solutions in [dargent2]). Mazzini [mazzini, mazzini2] explored the effect of the eclipsing constraint in averaging in more detail using a discontinuous model for the constraint. Tarzi [tarzi] incorporated solar radiation pressure into the model and introduced a multi-arc approach to averaging for bang-bang control, but did not account for the Earth’s shadow in an optimal way. Wu et al. [wu_avg] develop a method to analytically calculate the roots of the switching function in the averaging context and use a neural network for initial guess generation, but does not include eclipsing in an optimal manner or in examples. Many of these more recent averaging methods are compared by Petropolous [petropoulos_averaging]. The objective of this current work is to combine and extend the advances for the averaged dynamics case into a general and practical formulation for minimum fuel problems. The new model includes optimal handling of the eclipsing constraint, multi-arc averaging with novel application of the Leibniz integral rule, inclusion of generic dynamics perturbations, and variational equations for averaged dynamics for the first time.
@@
Dynamics averaging removes the fast periodic changes in the states and costates, leaving only the slowly changing secular motion. The resulting averaged trajectory will approximate the unaveraged trajectory to within , for integration times of up to [sanders2007averaging], where is typically a small parameter defined in terms of perturbation amplitude (control or dynamics). A numerical integrator can then take much larger time steps, which both reduces the numerical sensitivity to the initial costates and reduces computation time to solve the resulting system of ordinary differential equations. Dynamics averaging has also been used effectively in collocation methods that discretize and optimize the states and costates directly. Falck [falck2012] builds on the averaging work done by Kluever [kluever] to develop a general collocation method for the averaged minimum time problem. Olikara [olikara2018, olikara2019] was the first to introduce a collocation method for the averaged minimum fuel problem. Shannon [shannon2024] improves on Olikara’s method for optimal coasting in collocation and introduces a hybrid approach that switches to unaveraged dynamics in regimes where the averaging approximation deteriorates. The collocation approach allows for straightforward inclusion of path constraints, such as eclipsing, but typically requires the number of control variables to scale with time of flight.
@@
In practice, dynamics averaging is done by taking the Hamiltonian (with the embedded primer vector feedback law) and integrating over one period in time to find the averaged Hamiltonian. Derivatives needed for equations of motion (and the variational equations) are then calculated using the averaged Hamiltonian. Following the methods of Kechichian [kechichian] and Edelbaum et al. [edelbaum], the averaging integrals are approximated using a weighted sum with Gaussian Quadrature because they rarely can be calculated analytically. For the minimum fuel problem with bang-bang control, a significant number of quadrature points are required to accurately capture the thrusting and coasting proportions of the averaging period. The multi-arc averaging approach introduced by Tarzi [tarzi] for bang-bang control separates the integral over the Hamiltonian into different thrusting or ballistic “arcs”. The multi-arc approach reduces the number of quadrature points required to achieve a specific level of accuracy at the cost of root finding in each dynamics function call.
@@
One of the most important constraints for low-thrust trajectory design is the effect of the Earth’s shadow, which typically requires a spacecraft with an electric propulsion engine to stop thrusting. Incorporating this constraint in a way that maintains optimality is not simple in general, and especially difficult when using dynamics averaging. In a proper optimal control derivation, the eclipsing constraint is directly incorporated into the dynamics so that the control law is optimal with respect to the constraint [singh_eclipse, WU2021107146]. If, instead, the constraint is simply enforced in flight, using a thrust profile that did not consider eclipsing, then the control law is technically no longer optimal. Mazzini describes the importance of the dynamics-based control constraint with multiple examples, showing how it can provide significant fuel savings depending on the Sun position at launch [mazzini]. When the constraint is properly incorporated, the costates exhibit a discrete “jump” at eclipse entrance and exit. In averaged dynamics, these discrete jumps in a costate variable are not easily incorporated. Mazzini [mazzini] develops a method to directly calculate the discrete jumps by integrating approximate costate dynamics through a transitionary layer for eclipsing. The averaged costate jumps are then calculated by applying the averaging integral to costate equations of motion that include a Dirac delta function. The resulting form of the averaged costate jumps is sufficient for propagation of an averaged trajectory, but does not address calculation of the variational equations. The model introduced in this current work incorporates the costate jumps into the averaged dynamics using the Leibniz integral rule applied to multi-arc averaging. This new approach uses no approximations and enables straightforward calculation of the variational equations.
@@
In the current work, the model includes the multi-arc averaging approach introduced by Tarzi [tarzi] along with a novel application of the Leibniz integral rule to automatically incorporate the averaged costate jumps. The new approach is simple, efficient, optimal with respect to the eclipsing constraint, and allows for straight-forward calculation of the variational equations. The averaged costate jumps, however, approach a singularity as eclipsing arcs approach zero length. This singularity, also identified in Mazzini’s model, can significantly slow down or prevent propagation with a variable step integrator and result in accumulation of error. The new averaging model removes this singularity through a simple redefinition of the eclipsing constraint.
@@ 2 Optimal Control
Similar to prior works on this many-revolution low-thrust problem [junkinsTaheri, taheri, graham2016minimum], modified equinoctial orbital elements are used to reduce the number of singularities compared to Keplerian orbital elements [walker1986set]. These elements are defined in terms of the Keplerian orbital elements as:
where a is the semimajor axis, i is the inclination, e is the eccentricity, is the argument of perigee (AOP), is the right ascension of the ascxcending node (RAAN), and is the true anomaly (TA). Current time, , and overall time of flight, , are included in the state variables:
@@
In order to incorporate the effects of the Earth’s shadow directly into the dynamics, the conventional geometric “eclipsing function” is used [singh_eclipse]. The eclipsing function is a smooth function of the geocentric satellite position, the geocentric Sun position, the radius of the Sun, the radius of the Earth, and indirectly the state variable time. A conical shadow model is used, where the engine fully shuts off in umbra, penumbra, and antumbra (see Fig. 1)[williamsEclipse]. The eclipsing function is written as
@@
The spacecraft thrust is defined in the radial-transverse-normal (RTN) reference frame. This frame is defined by the spacecraft radial vector, the transverse velocity vector, and the normal vector that completes the orthogonal basis. The controllable variables are the thrust direction, , and a thrust modulating parameter, , with maximum and minimum thrusts given as and , respectively. The function is incorporated directly into the dynamics. The equations of motion are defined:
@@
where is an arbitrary perturbation, is the gravitation acceleration at Earth’s surface, and is the specific impulse. The matrices A, B contain the state dynamics equations for modified equinoctial orbital elements written in the RTN frame. They are derived by Kechichian [kechichian] and shown in the appendix. The arbitrary perturbation, , and the thrust unit vector, , are defined in the RTN frame, and must be a smooth function of state variables. This perturbation must be small enough to not significantly change the mean orbital elements over one period for the averaging assumptions to hold. Examples of generic perturbations that can be included are oblateness terms for the primary body, simple solar radiation pressure models, or simple drag models. In later examples, the perturbation is included.
@@
The minimum fuel cost function [bertrandepenoy] used in this model is defined as:
@@
where is the costate vector for the 6 orbital elements and two time related variables, and is the costate for the mass. The costate dynamics are then calculated as:
@@
To derive the optimal control law, the Hamiltonian is minimized with respect to the controls, thrust direction and thrust magnitude [kirk1970optimal]. For the direction, the thrust vector is anti-aligned with B [lawden_1963]:
The so-called "primer vector" [lawden_1963] is substituted back into Eq. (14) for an updated Hamiltonian:
The optimal is determined by the sign of the quantity multiplying it, the so-called "switching function":
This equation for the switching function can be further simplified. The quantities and are all greater than zero and will not affect the sign of Eq. (15). The variable will either be 1 and not affect the sign of Eq. (15) or it will be 0, in which case will have no effect on the equations of motion for the states or the costates. The latter case is technically a singular arc, but does not require further investigation because the singular control has no effect on the equations of motion. The optimal is thus determined by a simplified switching function:
which leads to so-called “bang-bang” control:
@@
This optimal is then substituted back into the Hamiltonian in Eq. (14). The derived optimal control law is now optimal with respect to the eclipsing constraint.
The traditional approach to maintaining optimality with a control constraint like eclipsing is to adjoin the Hamiltonian with a constraint function [brysonho, hull, clarke2013]. The adjoined Hamiltonian is then used to calculate the Weierstrass-Erdmann corner conditions, which determine how the costates “jump” at eclipse entry and exit:
where is a scalar Lagrange multiplier that can be calculated in terms of the states and costates. It should be noted that there is no discontinuous jump in the Hamiltonian due to the time regularization. Also, the primer vector, Eq. (LABEL:eq:primer_vector), and reduced switching function, Eq. (16), are continuous across eclipse entry and exit [wangtopputo]. Some recent works have successfully applied this traditional approach, also referred to as “interior point constraints”, to low-thrust trajectory optimization with eclipsing [wangtopputo, cerf2018, PONTANI2022549]. This traditional approach can be cumbersome, though, because it requires a trajectory propagation method that detects and stops exactly at every eclipse entry and exit. The eclipsing control constraint can also be incorporated into the dynamics with a smooth Heaviside approximation [FerrierEpenoy], which allows for straightforward trajectory propagation. The smooth, approximate Heaviside function approach will yield identical results to the discontinuous, true Heaviside function approach in the limit as smoothing parameters approach zero. The focus of this work is the averaged dynamics model, which does not exhibit discontinuous costate jumps, but does need to incorporate the averaged effect of these jumps into the averaged equations of motion. In the next section, the averaged dynamics are introduced, which use a novel application of the Leibniz integral rule to incorporate the effect of the costate jumps into the averaged equations of motion.
@@ 3.1 Basic method
The averaging approach used in this work follows that of Edelbaum et al. [edelbaum] and Kechichian [kechichian]. The process starts with the unaveraged Hamiltonian including the embedded optimal control. Next, the average is performed over one period in time, and this averaged Hamiltonian is used to calculate the equations of motion for the averaged states and costates. It should be noted that the optimal control does not need to be re-derived after averaging the Hamiltonian, because there are no direct constraints on the control [geffroy]. The averaging integral is converted from time to the fast moving true longitude, L. All other variables are held constant within the averaging integral. This process removes the fast periodic changes in the state and costate, leaving only the secular motion. Gaussian quadrature [quadrature] is used to approximate the integral and all other equations derived from it. The time averaged Hamiltonian, , over one period is defined here and rearranged slightly:
where a new term, , is defined for convenience and is only a function of the state variables, not the costates. The state and costate dynamics are calculated using the chain rule:
@@
where the weights, , and node points, , are calculated from Legendre polynomials [quadrature]. The dynamics equations are then calculated using quadrature:
The number of quadrature points required, n, to make the approximation sufficiently accurate depends on the complexity of the unaveraged dynamics within the given period of averaging. If there are sections of the given period with and without thrusting, then n must be large enough to accurately capture the proportion and relative locations of the two. This common [geffroy, mazzini], but brute force approach to calculating the averaged equations of motion can sometimes require such large n that the computational savings of averaged propagation are significantly reduced. Alternatively, a multi-arc approach improves the accuracy of approximation without requiring excessively large n.